• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Kelly, no RF

    Log entry time 15:59:12 on June25,2000

    Entry number 44897

    keyword=day shift summary




    Continue production running at thcm=0.

    0830: MCC adjusted pathlength
    0830: The rear x1 histogram has a ragged region of about 16 channels.
    1000: Rear x1 histogram is getting worse. Bad region about 32 channels
    1045: Most FPP spectra are now bad and reconstruction is affected.
    We've called Mark Jones about FPP and plant engineering
    about the temperature in the hadron hut. They report that the
    temperature reading is still about 78 deg, but perhaps it is
    worse near the electronics.
    1100: arranged for access at about 1130 with Mark and engineering
    1100: CHL alarm takes all RF off. MCC evaluating problem and
    recovery time.
    1145: enter Hall to check FPP multiplexing and hut air conditioning
    1315: leave hall. Air conditioner on rhs was off. Plant engineering
    changed filter and restarted unit. No obvious problem with
    signals, but timing shifts could be subtle. Running cosmics
    while hut cools down again.


    run events comments
    ==== ===== ========
    2534 1.3 M
    2535 1.3 M
    2536 1.3 M
    2537 0.75 M stopped because of CHL alarm
    2538 cosmics
    2539 cosmics, runtype hadron, ps1-5 ->1
    2540 cosmics, runtype twospect, ps1-5 -> 1
    2541 cosmics, ps4=1000, others 1.
    2542 cosmics, ps3=1, others=65535